# RoadBench: A Vision-Language Foundation Model and Benchmark for Road Damage Understanding

### v1

@@ 1 Introduction
Road infrastructure is the backbone of economic development and social connectivity, supporting efficient transportation, commerce, and access to essential services. However, damages such as cracks, potholes, and pavement deformation significantly reduce ride quality and, if left unaddressed, pose serious safety risks, resulting in substantial economic costs. To tackle these challenges, the research community has turned to computer vision to develop automated systems for road damage detection [2, 4, 83]. The emergence of several large-scale vision-only datasets (e.g., RDD2022 [4] with over 47,000 images of road surfaces across six countries) has catalyzed the development of deep learning models for pavement distress identification and classification [6, 52, 51].
@@
However, existing models are exclusively vision-based, relying solely on visual cues (e.g., image features), without considering textual descriptions of the damage [2, 4]. As a result, these models often struggle to capture fine-grained distinctions, such as the severity of a crack or its precise location, and may fail to generalize across diverse environmental and road conditions. Inspired by the remarkable success of multimodal approaches, like CLIP [50], which demonstrates a strong zero-shot recognition capability by learning joint image–text representations, we are motivated to explore whether multimodality could significantly improve accuracy, adaptability, and semantic understanding in real-world infrastructure monitoring scenarios. However, to our best knowledge, no publicly available dataset pairs road surface images with detailed textual descriptions, limiting the development of such multimodal solutions in this domain. Moreover, well-known large foundation models (e.g., GPT-4o [48] and DeepSeek-VL [43]) yield poor performance on road damage understanding (see Fig 6 for details), a challenge often tied to the transferability of general-purpose representations to specialized tasks [46, 87, 86].
@@
To fill this gap, we introduce RoadBench, the first multimodal benchmark for road damage understanding, along with RoadCLIP, the first vision–language model tailored for joint image–text learning in this domain, built on this dataset. RoadBench includes 100,000 high-resolution road images, with each paired with a detailed and reliable description of the scene’s pavement condition, generated by a state-of-the-art generative language model (i.e., GPT-4o [48]). The descriptions are employed to reflect real-world conditions, including diverse environments (e.g., urban vs. rural), varied weather and lighting, and a wide range of damage types and severities. We show several examples in Figure 1. RoadCLIP comprises two key modules: a disease-aware positional encoding module and a domain-specific prior injection mechanism. The disease-aware positional encoding module injects knowledge of road geometry and common damage localization patterns into the visual branch of the multimodal backbone, making the learned representations sensitive to the road surface, where a damage occurs (e.g., wheel-path cracks and shoulder cracks), and the spatial scope of the damage, well aligned with human inspectors’ professional actions of checking both location and spread of a damage. The domain-specific prior injection equips the model with expert knowledge about typical “road diseases” (e.g., the co-occurrence of certain crack patterns or textural cues indicating material fatigue). This prior is integrated during training for image–text alignment, acting as guidance that makes multimodal representations more discriminative for our task. With the two primary innovations, RoadCLIP learns a joint embedding space where images and descriptions of road damage are tightly aligned, facilitating more accurate cross-modal understanding.
@@ 2 Related Work
Dataset Modality Samples Resolution Disease Types Task Type Availability
RDD2022 [4] Vision-only 47,420 Variable Crack, Pothole Detection Public
TD-RD [70] Vision-only 7,088 3840×2160 Crack, Pothole, Repair Detection Public
CRDDC’22 [3] Vision-only 11,720 Variable Crack, Patch, Spalling Detection Public
CNRDD [83] Vision-only 9,053 Variable Crack, Patch, Rutting Detection Public
GAPs384 [8] Vision-only 1,969 1920×1080 Crack, Pothole Detection Public
CFD [56] Vision-only 1,180 480×320 Crack Detection Public
RoadBench (Ours) Vision+Text (Multimodal) 100,000 3840×2160 Ten types in all, more details see Fig3 Multimodal Detection, Retrieval, Captioning, QA Public
@@ 2.1 Road Damage Detection Benchmarks
Existing road damage datasets and benchmarks are predominantly vision-only, focusing on detecting and localizing road surface defects from images [3, 2]. The Road Damage Dataset 2020 (RDD2020) [2] contains over 26,000 road images from multiple countries with more than 31,000 annotated damage instances, while the recent RDD2022 dataset [4] expanded this to 47,420 images across six countries, featuring over 55,000 labeled instances. Released as part of the CRDDC 2022 competition [3], these datasets have spurred the development of robust detection models using advanced architectures like YOLO series [51, 6, 78] and Transformer-based [7, 93, 90, 18, 30, 88]. Other specialized datasets include TD-RD [69], which offers aerial-view images from drone photography, GAPs384 [8] and CFD [56] for specific damage types, and CNRDD [83] featuring Chinese road conditions. However, the critical limitation across all these datasets is their unimodal nature—they contain only images and annotations without any natural language descriptions, which restricts contextual understanding, generalization capabilities [87], and semantic reasoning about damage characteristics [55, 57, 62, 34, 16, 84, 80, 35, 64, 68, 20, 67]. Detailed information on these datasets is presented in Table 1.
@@ 2.2 Vision-Language Models
There have been rapid advances in vision-language models that jointly learn from images and text [37, 29, 21, 26, 24, 14, 45, 33, 11, 54, 10, 91, 22, 63, 75, 23, 36]. Contrastive models like CLIP [50] align visual and textual representations using large-scale image–text pairs, enabling strong zero-shot performance. Unified architectures such as BLIP [31] and Florence [82] extend these capabilities to both understanding and generation through web-scale multimodal pretraining. Recent advances have further integrated powerful language models with visual processing [28, 5, 42, 38, 22, 25, 58, 65, 77, 85], including techniques like test-time adaptation [86] and prompt tuning [71, 72], as seen in models like LLaVA [40], GPT-4V [48], and Flamingo [1], which support sophisticated multimodal reasoning and contextual understanding. Evaluation typically uses benchmarks such as MS COCO Captions [39] and Flickr30k [49] for captioning and retrieval, VQAv2 [19] for visual question answering, RefCOCO [81] for referring expressions, and ScienceQA [44] for scientific reasoning. Despite this progress, a significant gap exists in domain-specific multimodal benchmarks. While similar efforts are emerging in other fields like medicine [73], none of the existing evaluation frameworks target specialized applications like road damage detection or infrastructure assessment, limiting our understanding of how these powerful models perform in critical real-world domains where both visual recognition and linguistic description are essential. .
@@ 3.1 Data Curation
Data Generation. We firstly consult domain experts to establish a standardized vocabulary including road disease types, environmental conditions, and beyond. This expert-informed terminology guides the construction of textual prompts, which are subsequently utilized by the multimodal GPT-4o model to synthesize high-resolution (3840×2160 pixels) road surface images accompanied by corresponding textual descriptions. The generated outputs are subsequently subjected to human expert verification to ensure accuracy and relevance, a practice also found effective in other domains like medical image synthesis [61]. A mixed-method labeling approach is then applied to enhance label quality, followed by a final composition step that integrates the image, caption, and corresponding label into the RoadBench dataset.
@@ 3.3 Is Synthetic Data Reliable?
Recent advances in data-driven models have highlighted the importance of large, diverse, and high-quality datasets. However, collecting and annotating real-world data is often costly, time-consuming, and privacy-sensitive, particularly in domains such as autonomous driving or road damage assessment. Consequently, synthetic data generation has emerged as a viable alternative [61, 12], enabling controlled sampling, scalable labeling, and simulation of rare or hazardous scenarios. The reliability of synthetic data has been studied across various vision tasks. For example, Richter et al. [53] demonstrated that synthetic datasets generated from video game engines (e.g., GTA V) can effectively train semantic segmentation models. Similarly, Tremblay et al. [59] showed that models trained on synthetic objects can generalize well to real-world object detection. More recently, synthetic datasets like Synscapes [66] and CARLA [15] have been widely adopted in autonomous driving research due to their realism and label precision. In this work, synthetic data serves a complementary role to real-world annotations. By leveraging a simulation pipeline augmented with text-guided prompt generation and selective augmentation, we ensure that the synthesized samples retain semantic fidelity and style diversity while remaining free of privacy concerns such as faces or license plates.
@@ 4.1 Overall Architecture
RoadCLIP adopts a dual-encoder architecture building upon the CLIP framework [50], as shown in Figure 4. It consists of a vision encoder $f(\cdot)$ and a text encoder $g(\cdot)$ that project images and textual descriptions, respectively, into a shared $d$-dimensional latent space. In the vision module, we use a Transformer-based image encoder (as in CLIP, e.g. a ViT) that processes the input image as a sequence of patch embeddings. The text module is a Transformer language model encoding the input description or label. Both encoders produce a feature vector of dimension $d$ (using the Transformer’s [CLS] token or global pooled representation), which is then $\ell_{2}$-normalized and passed through a linear projection (also of dimension $d$) to produce the final image embedding $\mathbf{z}_{i}$ and text embedding $\mathbf{z}_{t}$. To adapt RoadCLIP for road-damage-aware representation learning, we incorporate two novel modules into its architecture: a Disease-aware Positional Encoding (DaPE) module integrated into the vision encoder, and a Domain-Specific Prior Injection mechanism affecting both the image and text representations.
@@ 4.4 Training Objective
Image-Text Contrastive Loss. The primary objective is a bidirectional InfoNCE loss [47] encouraging high similarity between matched image–text pairs $(I_{i},T_{i})$ and low similarity for mismatches. Let $\mathbf{z}_{i}=f(I_{i})$ and $\mathbf{z}_{i}^{+}=g(T_{i})$ denote the embeddings from the image and text encoders, respectively. The loss is:
@@
Single-modal Vision-only Methods Multimodal Vision-Language Models
Method Modality SLA(%) Method Modality ZS Acc.(%) Recall@1(%) Recall@5(%) Recall@10(%) SLA(%)
YOLOv10-n [NeurIPS24] [60] Vision 46.5 CLIP [ICML21] [50] Vision+Text 63.8 40.3 63.2 71.4 41.5
YOLOv10-s [NeurIPS24] [60] Vision 47.3 BLIP-2 [ICML23] [32] Vision+Text 67.9 43.7 66.5 74.0 46.3
YOLOS-ti [NeurIPS21] [17] Vision 45.3 LLaVA [NeurIPS23] [40] Vision+Text 69.1 46.0 68.9 76.2 48.0
YOLOS-s [NeurIPS21] [17] Vision 46.8 MiniGPT-4 [arXiv23] [92] Vision+Text 68.7 44.9 67.5 75.3 47.1
PP-PicoDet [arXiv21] [79] Vision 47.0 InstructBLIP [NeurIPS23] [13] Vision+Text 70.2 47.6 70.1 77.8 49.2
RT-DETR [CVPR24] [89] Vision 48.6 mPLUG-Owl [arXiv23] [76] Vision+Text 71.0 49.3 71.8 78.6 50.7
Lite-DETR [CVPR23] [30] Vision 45.3 Otter [arXiv23] [27] Vision+Text 71.4 49.7 72.2 79.1 50.9
FR-CNN [NeurIPS15] [52] Vision 39.9 GPT-4V [arXiv23] [48] Vision+Text 74.5 52.6 74.0 80.8 55.4
SSD-VGG16 [ECCV16] [41] Vision 38.7 DeepSeek-VL [arXiv24] [43] Vision+Text 72.9 50.1 72.3 79.5 53.0
TD-YOLOV10 [ICASSP25] [70] Vision 49.0 RoadCLIP (Ours) Vision+Text 78.6 58.4 78.1 84.6 61.9
@@ 5.1 Experimental Setup
Datasets. We evaluate model performance primarily on our RoadBench dataset and further evaluate cross-dataset generalization on three widely used visual-only datasets: TD-RD [69], CNRDD [83], and CRDDC’22 [3].
Evaluation Metrics. Zero-Shot Classification Accuracy (ZS Acc.) measures the model’s ability to classify road damage categories without seeing labeled training data for those specific types, reflecting semantic alignment and domain adaptation capability. Image-Text Retrieval Accuracy is reported as Recall@$k$ (with $k=1,5,10$), evaluating how accurately the model retrieves the correct caption given an image (or vice versa), a standard protocol in vision-language literature [50, 32]. Additionally, we introduce Semantic Localization Accuracy (SLA) to quantify the model’s ability to correctly ground and localize defects in the image guided by textual descriptions—a task inspired by recent advances in referring expression grounding [74]. SLA is computed by measuring IoU overlap between predicted and annotated binary masks generated via text-guided attention.
Implementation Details. We firstly initialize our model with CLIP [50] pretraining weights and then fine-tune on RoadBench using a contrastive InfoNCE-style objective. Unless otherwise stated, the batch size is set to 128, learning rate to $1e^{-4}$, and training runs for 20 epochs.

### v3

@@ 1 Introduction
Road infrastructure is the backbone of economic development and social connectivity, supporting efficient transportation, commerce, and access to essential services. However, damages such as cracks, potholes, and pavement deformation significantly reduce ride quality and, if left unaddressed, pose serious safety risks, resulting in substantial economic costs. To tackle these challenges, the research community has turned to computer vision to develop automated systems for road damage detection [arya2021rdd2020, arya2024rdd2022, zhang2022new]. The emergence of several large-scale vision-only datasets (e.g., RDD2022 [arya2024rdd2022] with over 47,000 images of road surfaces across six countries) has catalyzed the development of deep learning models for pavement distress identification and classification [bochkovskiy2020yolov4, ren2015faster, redmon2016you].
@@
However, existing models are exclusively vision-based, relying solely on visual cues (e.g., image features), without considering textual descriptions of the damage [arya2021rdd2020, arya2024rdd2022]. As a result, these models often struggle to capture fine-grained distinctions, such as the severity of a crack or its precise location, and may fail to generalize across diverse environmental and road conditions. Inspired by the remarkable success of multimodal approaches, like CLIP [radford2021learning], which demonstrates a strong zero-shot recognition capability by learning joint image–text representations, we are motivated to explore whether multimodality could significantly improve accuracy, adaptability, and semantic understanding in real-world infrastructure monitoring scenarios. However, to our best knowledge, no publicly available dataset pairs road surface images with detailed textual descriptions, limiting the development of such multimodal solutions in this domain. Moreover, well-known large foundation models (e.g., GPT-4o [openai2023gpt4] and DeepSeek-VL [lu2024deepseekvl]) yield poor performance on road damage understanding (see Fig 6 for details), a challenge often tied to the transferability of general-purpose representations to specialized tasks [NEURIPS2024_d3602fc9, zhang2025dpcore, Zhang_2025_WACV].
@@
To fill this gap, we introduce RoadBench, the first multimodal benchmark for road damage understanding, along with RoadCLIP, the first vision–language model tailored for joint image–text learning in this domain, built on this dataset. RoadBench includes 100,000 high-resolution road images, with each paired with a detailed and reliable description of the scene’s pavement condition, generated by a state-of-the-art generative language model (i.e., GPT-4o [openai2023gpt4]). The descriptions are employed to reflect real-world conditions, including diverse environments (e.g., urban vs. rural), varied weather and lighting, and a wide range of damage types and severities. We show several examples in Figure 1. RoadCLIP comprises two key modules: a disease-aware positional encoding module and a domain-specific prior injection mechanism. The disease-aware positional encoding module injects knowledge of road geometry and common damage localization patterns into the visual branch of the multimodal backbone, making the learned representations sensitive to the road surface, where a damage occurs (e.g., wheel-path cracks and shoulder cracks), and the spatial scope of the damage, well aligned with human inspectors’ professional actions of checking both location and spread of a damage. The domain-specific prior injection equips the model with expert knowledge about typical “road diseases” (e.g., the co-occurrence of certain crack patterns or textural cues indicating material fatigue). This prior is integrated during training for image–text alignment, acting as guidance that makes multimodal representations more discriminative for our task. With the two primary innovations, RoadCLIP learns a joint embedding space where images and descriptions of road damage are tightly aligned, facilitating more accurate cross-modal understanding.
@@ 2 Related Work
[…]
RDD2022 [arya2024rdd2022] Vision-only 47,420 Variable Crack, Pothole Detection Public
TD-RD [xiao2025tdrdtopdownbenchmarkrealtime] Vision-only 7,088 3840×2160 Crack, Pothole, Repair Detection Public
CRDDC’22 [arya2022crowdsensing] Vision-only 11,720 Variable Crack, Patch, Spalling Detection Public
CNRDD [zhang2022new] Vision-only 9,053 Variable Crack, Patch, Rutting Detection Public
GAPs384 [cha2024deep] Vision-only 1,969 1920×1080 Crack, Pothole Detection Public
CFD [shi2016automatic] Vision-only 1,180 480×320 Crack Detection Public
[…]
@@ 2.1 Road Damage Detection Benchmarks
Existing road damage datasets and benchmarks are predominantly vision-only, focusing on detecting and localizing road surface defects from images [arya2022crowdsensing, arya2021rdd2020]. The Road Damage Dataset 2020 (RDD2020) [arya2021rdd2020] contains over 26,000 road images from multiple countries with more than 31,000 annotated damage instances, while the recent RDD2022 dataset [arya2024rdd2022] expanded this to 47,420 images across six countries, featuring over 55,000 labeled instances. Released as part of the CRDDC 2022 competition [arya2022crowdsensing], these datasets have spurred the development of robust detection models using advanced architectures like YOLO series [redmon2016you, bochkovskiy2020yolov4, yu2021pp] and Transformer-based [carion2020end, zhu2021deformable, zhao2024detrs, fang2021you, li2023lite, zhang2021vit]. Other specialized datasets include TD-RD [xiao2025tdrd], which offers aerial-view images from drone photography, GAPs384 [cha2024deep] and CFD [shi2016automatic] for specific damage types, and CNRDD [zhang2022new] featuring Chinese road conditions. However, the critical limitation across all these datasets is their unimodal nature—they contain only images and annotations without any natural language descriptions, which restricts contextual understanding, generalization capabilities [zhang2025dpcore], and semantic reasoning about damage characteristics [sharma2018conceptual, subramanian2020medicat, wang2024ensemble, li2024yolox, dugalam2024development, zhang2024real, yu2024road, 9904187, wang2024drivedreamer, wu2024seesr, he2024can, wu2024one]. Detailed information on these datasets is presented in Table 1.
@@ 2.2 Vision-Language Models
There have been rapid advances in vision-language models that jointly learn from images and text [li2025benchmark, li2022elevater, hong2020more, kuckreja2024geochat, jain2024semantic, dong2024changeclip, marino2019ok, li2024survey, chen2023benchmarking, roberts2024image2struct, chen2024viseval, zheng2022vlmbench, hong2024cogagent, wang2024cogvlm, yao2024tcp, huang2024adapting, li2024cycleyoloefficientrobustframework]. Contrastive models like CLIP [radford2021learning] align visual and textual representations using large-scale image–text pairs, enabling strong zero-shot performance. Unified architectures such as BLIP [li2022blip] and Florence [yuan2021florence] extend these capabilities to both understanding and generation through web-scale multimodal pretraining. Recent advances have further integrated powerful language models with visual processing [li2024seed, bitton2023visit, liu2024ii, lin2024vila, hong2024cogagent, kayser2021vil, tankala2024wikido, wang2024journeybench, yin2023lamm, zhang2024unveiling], including techniques like test-time adaptation [Zhang_2025_WACV] and prompt tuning [xiao2025visualinstanceawareprompttuning, xiao2025visualvariationalautoencoderprompt], as seen in models like LLaVA [liu2023llava], GPT-4V [openai2023gpt4], and Flamingo [alayrac2022flamingo], which support sophisticated multimodal reasoning and contextual understanding. Evaluation typically uses benchmarks such as MS COCO Captions [lin2014microsoft] and Flickr30k [plummer2015flickr30k] for captioning and retrieval, VQAv2 [goyal2017making] for visual question answering, RefCOCO [yu2016modeling] for referring expressions, and ScienceQA [lu2022learn] for scientific reasoning. Despite this progress, a significant gap exists in domain-specific multimodal benchmarks. While similar efforts are emerging in other fields like medicine [xiao2025medicalimages], none of the existing evaluation frameworks target specialized applications like road damage detection or infrastructure assessment, limiting our understanding of how these powerful models perform in critical real-world domains where both visual recognition and linguistic description are essential. .
@@ 3.1 Data Curation
Data Generation. We firstly consult domain experts to establish a standardized vocabulary including road disease types, environmental conditions, and beyond. This expert-informed terminology guides the construction of textual prompts, which are subsequently utilized by the multimodal GPT-4o model to synthesize high-resolution (3840×2160 pixels) road surface images accompanied by corresponding textual descriptions. The generated outputs are subsequently subjected to human expert verification to ensure accuracy and relevance, a practice also found effective in other domains like medical image synthesis [wang2025doctor]. A mixed-method labeling approach is then applied to enhance label quality, followed by a final composition step that integrates the image, caption, and corresponding label into the RoadBench dataset.
@@ 3.3 Is Synthetic Data Reliable?
Recent advances in data-driven models have highlighted the importance of large, diverse, and high-quality datasets. However, collecting and annotating real-world data is often costly, time-consuming, and privacy-sensitive, particularly in domains such as autonomous driving or road damage assessment. Consequently, synthetic data generation has emerged as a viable alternative [wang2025doctor, chung2025soksyntheticimagesreplace], enabling controlled sampling, scalable labeling, and simulation of rare or hazardous scenarios. The reliability of synthetic data has been studied across various vision tasks. For example, Richter et al. [richter2016playing] demonstrated that synthetic datasets generated from video game engines (e.g., GTA V) can effectively train semantic segmentation models. Similarly, Tremblay et al. [tremblay2018training] showed that models trained on synthetic objects can generalize well to real-world object detection. More recently, synthetic datasets like Synscapes [wrenninge2018synscapes] and CARLA [dosovitskiy2017carla] have been widely adopted in autonomous driving research due to their realism and label precision. In this work, synthetic data serves a complementary role to real-world annotations. By leveraging a simulation pipeline augmented with text-guided prompt generation and selective augmentation, we ensure that the synthesized samples retain semantic fidelity and style diversity while remaining free of privacy concerns such as faces or license plates.
@@ 4.1 Overall Architecture
RoadCLIP adopts a dual-encoder architecture building upon the CLIP framework [radford2021learning], as shown in Figure 4. It consists of a vision encoder $f(\cdot)$ and a text encoder $g(\cdot)$ that project images and textual descriptions, respectively, into a shared $d$-dimensional latent space. In the vision module, we use a Transformer-based image encoder (as in CLIP, e.g. a ViT) that processes the input image as a sequence of patch embeddings. The text module is a Transformer language model encoding the input description or label. Both encoders produce a feature vector of dimension $d$ (using the Transformer’s [CLS] token or global pooled representation), which is then $\ell_{2}$-normalized and passed through a linear projection (also of dimension $d$) to produce the final image embedding $\mathbf{z}_{i}$ and text embedding $\mathbf{z}_{t}$. To adapt RoadCLIP for road-damage-aware representation learning, we incorporate two novel modules into its architecture: a Disease-aware Positional Encoding (DaPE) module integrated into the vision encoder, and a Domain-Specific Prior Injection mechanism affecting both the image and text representations.
@@ 4.4 Training Objective
Image-Text Contrastive Loss. The primary objective is a bidirectional InfoNCE loss [oord2018representation] encouraging high similarity between matched image–text pairs $(I_{i},T_{i})$ and low similarity for mismatches. Let $\mathbf{z}_{i}=f(I_{i})$ and $\mathbf{z}_{i}^{+}=g(T_{i})$ denote the embeddings from the image and text encoders, respectively. The loss is:
@@
[…]
YOLOv10-s [NeurIPS24] [wang2024yolov10] Vision 47.3 BLIP-2 [ICML23] [li2023blip] Vision+Text 67.9 43.7 66.5 74.0 46.3
YOLOS-ti [NeurIPS21] [Fang2021YouOL] Vision 45.3 LLaVA [NeurIPS23] [liu2023llava] Vision+Text 69.1 46.0 68.9 76.2 48.0
YOLOS-s [NeurIPS21] [Fang2021YouOL] Vision 46.8 MiniGPT-4 [arXiv23] [zhu2023minigpt] Vision+Text 68.7 44.9 67.5 75.3 47.1
PP-PicoDet [arXiv21] [yu2021pppicodetbetterrealtimeobject] Vision 47.0 InstructBLIP [NeurIPS23] [dai2023instructblip] Vision+Text 70.2 47.6 70.1 77.8 49.2
RT-DETR [CVPR24] [Zhao_2024_CVPR] Vision 48.6 mPLUG-Owl [arXiv23] [ye2023mplug] Vision+Text 71.0 49.3 71.8 78.6 50.7
Lite-DETR [CVPR23] [li2023lite] Vision 45.3 Otter [arXiv23] [li2023ottermultimodalmodelincontext] Vision+Text 71.4 49.7 72.2 79.1 50.9
FR-CNN [NeurIPS15] [ren2015faster] Vision 39.9 GPT-4V [arXiv23] [openai2023gpt4] Vision+Text 74.5 52.6 74.0 80.8 55.4
SSD-VGG16 [ECCV16] [liu2016ssd] Vision 38.7 DeepSeek-VL [arXiv24] [lu2024deepseekvl] Vision+Text 72.9 50.1 72.3 79.5 53.0
TD-YOLOV10 [ICASSP25] [xiao2025tdrdtopdownbenchmarkrealtime] Vision 49.0 RoadCLIP (Ours) Vision+Text 78.6 58.4 78.1 84.6 61.9
@@ 5.1 Experimental Setup
Datasets. We evaluate model performance primarily on our RoadBench dataset and further evaluate cross-dataset generalization on three widely used visual-only datasets: TD-RD [xiao2025tdrd], CNRDD [zhang2022new], and CRDDC’22 [arya2022crowdsensing].
Evaluation Metrics. Zero-Shot Classification Accuracy (ZS Acc.) measures the model’s ability to classify road damage categories without seeing labeled training data for those specific types, reflecting semantic alignment and domain adaptation capability. Image-Text Retrieval Accuracy is reported as Recall@$k$ (with $k=1,5,10$), evaluating how accurately the model retrieves the correct caption given an image (or vice versa), a standard protocol in vision-language literature [radford2021learning, li2023blip]. Additionally, we introduce Semantic Localization Accuracy (SLA) to quantify the model’s ability to correctly ground and localize defects in the image guided by textual descriptions—a task inspired by recent advances in referring expression grounding [yang2023semantics]. SLA is computed by measuring IoU overlap between predicted and annotated binary masks generated via text-guided attention.
Implementation Details. We firstly initialize our model with CLIP [radford2021learning] pretraining weights and then fine-tune on RoadBench using a contrastive InfoNCE-style objective. Unless otherwise stated, the batch size is set to 128, learning rate to $1e^{-4}$, and training runs for 20 epochs.